# 20xx Vol. xx No. x, XXX

### v1

@@ Abstract
With a growing sample of fast radio bursts (FRBs), we investigate the energy budget of different power sources within the framework of magnetar starquake triggering mechanism. During a starquake, the energy can be released in any form through magnetic, strain, rotational, and gravitational energies. Following findings are revealed: 1. The crust can store a free magnetic energy of the amount of at least $6.3\times 10^{46}$ erg via toroidal fields, with frequent starquakes happening due to the instability of the crust. 2. The strain energy develops as a rigid object spins down, which can be released during a global starquake accompanied by a glitch. However, it takes a long time to accumulate enough strain energy via spin-down. 3. The rotational energy of a magnetar with $P\lesssim 0.1\rm\,s$ can match the energy and luminosity budget of FRBs. 4. The budget of the total gravitational energy is high, but the mechanism and efficiency of converting this energy to radiation deserve further exploration.
@@ 1 Introduction
Fast radio bursts (FRBs) are millisecond-duration flashes, signifying extrem coherent radiation due to their extremely high bright temperatures. Although they are similar in some aspects to single pulses from radio pulsars, the energy budget of FRBs is much higher. As a result, the physical origin(s) of FRBs is still a mystery (Zhang 2023).
@@
A strong connection between magnetars and FRBs was established since the discovery of FRB 20200428D, an FRB-like burst from a Galactic magnetar, SGR J1935+2154 (Bochenek et al. 2020; CHIME/FRB Collaboration et al. 2020). Based on the distance from the emission site to the central engine, magnetar FRB models can be generally divided into two categories: emission within the magnetosphere (pulsar-like models, Katz 2014; Kumar et al. 2017; Yang & Zhang 2018; Wang et al. 2019; Lu et al. 2020; Cooper & Wijers 2021; Wang et al. 2022; Zhang 2022; Liu et al. 2023; Wang et al. 2023b; Qu & Zhang 2024), and emission from a relativistic shock region far outside the magnetosphere (GRB-like models, Metzger et al. 2019; Beloborodov 2020; Margalit et al. 2020; Chen et al. 2022; Khangulyan et al. 2022). The observations of polarization measurements, including diverse polarization angle swings (Luo et al. 2020b) and high degrees of circular polarization (Jiang et al. 2022; Xu et al. 2022; Zhang et al. 2023b), favor the pulsar-like models (Qu & Zhang 2023), even though there are also some attempts to interpret these phenomena within the GRB-like models (Iwamoto et al. 2024).
@@
Starquakes from magnetars have been proposed as a promising triggering mechanism for FRBs (Wang et al. 2018; Suvorov & Kokkotas 2019; Wadiasingh & Timokhin 2019; Jiang et al. 2020; Li et al. 2022a). The FRB burst sequences share remarkably similar characteristics in energy distribution and temporal occurrence with Earthquakes (Du et al. 2023; Totani & Tsuzuki 2023; Wang et al. 2023a; Tsuzuki et al. 2024). The quake model predicts that FRBs are associated with glitches and quasi-periodic oscillations (QPOs). Surprisingly, a giant glitch was measured days before FRB 20200428D (Mereghetti et al. 2020; Li et al. 2021a; Ridnaia et al. 2021), and later another glitch was found accompanied by three FRB-like radio bursts in subsequent days (Younes et al. 2023). A QPO of 40 Hz was reported from an X-ray burst during the active epoch of FRB 20200428D (Li et al. 2022b).
@@
Several FRB sources have been observed to repeat more than a thousand of times (Li et al. 2021b; Xu et al. 2022; Zhang et al. 2023b), suggesting that an active central engine and a large energy reservoir are needed. The prospect of detecting more than 10 thousand bursts from an FRB source would pose significant challenges on the potential energy budget of a magnetar. In this paper, we investigate the energy budget and the mechanism that can be released, within the framework of starquake model for a neutron star or a magnetar. We consider the energy budget of repeating FRBs in Section 2. We discuss the energy release during starquakes by invoking magnetic, strain, rotational, and gravitational energy in Section 3. The results are discussed in Section 4 and summarized in Section 5.
@@ 2 Energy budget of an FRB source
As a cosmological burst, an FRB can release a huge amount of energy. For a nominal Gpc distance $D$, the burst energy can be calculated as
$\delta E$ $=4\pi D^{2}f_{b}F_{\nu}\Delta\nu$ (1)
@@
where $f_{b}=\delta\Omega/(4\pi)$ is beam factor for the individual burst, $F_{\nu}$ is the received fluence and $\Delta\nu$ is the bandwidth. The emission solid angle $\delta\Omega$ could be as small as $1/(4\gamma^{2})\sim 2.5\times 10^{-5}$ due to relativistic beaming (Wang et al. 2022; Zhang 2022). Although some repeaters have been reported to emit more than thousands of bursts (Li et al. 2021b; Xu et al. 2022; Zhou et al. 2022), there are still a significant number of bursts that remain unobserved due to the beaming effect. There might be a global emission beam factor $F_{b}=\Delta\Omega/(4\pi)$ that is larger than that of individual bursts (Zhang 2023). When considering the total burst energy/luminosity, the global beam factor should be adopted rather than the beam factor of an individual burst.
It is necessary to introduce the luminosity/energy function of FRBs when a large sample of FRBs is considered. Some statistical works show that the luminosity/energy function of FRBs can be characterized by a power law distribution, $N(E)\propto E^{-p}$, with indices of $1-2$ (Luo et al. 2018; Lu & Piro 2019; Lu et al. 2020; Luo et al. 2020a). If the index is not strictly equal to 1 or 2, the average burst energy is calculated to
@@
Note that most bursts are missing due to the duty cycle. The total observed number and their energy budget of some actively repeating FRBs without duty cycle are summarized in Table 1. The physical duty cycle and burst rate evolution are complex. We assume that the repeater source has a lifetime $\tau_{\rm life}$ and consider the average bursting rate to be independent of energy. If the activity level remains unchanged during the lifetime, the total energy budget of the source over the lifefime can be estimated as (Zhang 2023)
@@
20121102A 2370 $4.1\times 10^{41}$ Spitler et al. (2016); Michilli et al. (2018); Zhang et al. (2018);
@@ 3 Energy release during starquake
Starquakes as a leading scenario to trigger FRBs can suddenly release a huge amount of energy. During the crustal motion, an Alfvén wave packet can be launched from the stellar surface. This process can extract and release magnetic energy from the crust (and possibly even from the interior of the crust) into the magnetosphere. Some strain energy may also be released when it is accumulated large enough. The star may release gravitational energy when its volume shrinks abruptly, especially in the case of accretion which can cause a substantial mass and gravity gain. Furthermore, a rotating magnetized neutron star can continuously convert its rotational energy into radiation from the magnetosphere. We discuss the release mechanisms of these energies independently in the following.
@@ 3.1 Magnetic energy released of starquake
Within the magnetar scenario, the magnetic energy can be released from the crust into the magnetosphere when the pressure induced by the internal magnetic field exceeds a threshold stress. The magnetic energy dominates the crustal deformation energy if the magnetic field strength $B>B_{\mu}=(4\pi\mu)^{1/2}=5.6\times 10^{15}\,\rm G$, where $\mu\approx 2.5\times 10^{30}\,\rm erg\,cm^{-3}$ is the shear modulus of the crust (Thompson & Duncan 1995). Suppose the magnetic field is deformed away from magnetostatic equilibrium by an amount of $\delta B$. In that case, the elastic stress suddenly balances the Maxwell stress, i.e., $\mu\epsilon\simeq B\delta B/(4\pi)$, where $\epsilon$ is the strain of the crust. The largest yield strain of the crust is $\sim 10^{-4}-10^{-2}$. The effect of ohmic decay is neglected for young magnetars.
@@
where $l$ is the patch size for the magnetic energy release, $B_{s}$ is the magnetic field strength on the stellar surface and $\Delta R$ is the crustal thickness. There is a height where charge density is insufficient to supply the current required by the wave and an electric field is formed to accelerate particle bunches, converting the magnetic energy into emissions. Even with the correction of the efficiency factor $\eta=10^{-4}$, the energy release can sustain a burst emitted by ultra-relativistic charges.
The timescale of the energy release process in an area of $l^{2}$ is related to some instability growth. A displacement of the magnetic footpoints on the surface can be caused by the diffusion of the internal magnetic field, which is reminiscent of the turbulent convective motions in the Sun. The exchange in the positions of the footpoints is driven by the interchange instability (Thompson & Duncan 1993). We take the density of the neutron drip $\rho_{n}$ as the crustal density and then the Alfvén speed in the crust is written as $B/(4\pi\rho_{n})^{1/2}$. The growth time of the instability driven by the external magnetic field is
$t_{\rm ex}\sim l/v_{A}\lesssim 2.2\times 10^{-5}\,{\rm s}\left(\frac{l}{10^{4}\,\rm cm}\right)\left(\frac{B}{10^{15}\,\rm G}\right)^{-1}.$ (5)
@@
The solid crust of a magnetar plays a crucial role in various high-energy activities such as FRB. Elastic stress accumulates with the evolution of the internal $B$-field and stores a large amount of energy in the form of elastic stress. Starquakes can occur anywhere on the star’s surface in principle. The total magnetic energy stored in the crust is
@@
where $R$ is the stellar radius. The burst number supported by the crustal magnetic energy is estimated by
@@
The magnetic energy stored in the magnetosphere is higher than that stored in the crust, and can support $\sim 10^{5}$ bursts. However, it is difficult to release such magnetic energy in a dipolar configuration, because the dipole is the ground state of the magnetic field. The free energy that magnetic fields can release favors multipoles or twisted magnetic fields.
Some toroidal magnetic field components might anchor in a highly conducting crust of a neutron star (Thompson et al. 2002). The toroidal magnetic field can be transferred from the interior to the crust via crustal differential rotation along gravitational equipotential surfaces. A sudden crust quake can twist the magnetic field in the outer magnetosphere (Beloborodov 2009). The magnetic energy of the twisted field is then launched into the magnetosphere. If the toroidal magnetic field in the crust is $10^{16}$ G, the energy can support $\sim 10^{6}$ bursts based on Equation (8). The elastic strain can hardly balance the magnetic strain because of $B>B_{\mu}$. Therefore, frequent starquakes may occur at the surface. A simulation has found that if the magnetic field has a strong toroidal component, the crust is most prone to rupture, and in this case, the epicenter of crustal earthquakes is near the equator (Lander et al. 2015). By assuming that the magnetic strain is comparable to gravity, there might be an upper limit of magnetic field of $\sim 10^{18}$ G, leading to $\sim 10^{10}$ bursts.
3.2 Strain energy during global quakes
During the crack motion, the elastic energy released in the crust excites oscillations, and the induced electric field accelerates charges to stream outflow (Tsygan 1975; Fabian et al. 1976; Muslimov & Tsygan 1986; Epstein 1988; Blaes et al. 1989). The strain energy released in the crack region of $l^{2}$ is given by
$\delta E_{\rm str}$ $=2\pi l^{2}\Delta R\mu\epsilon^{2}$ (11)
$=4.7\times 10^{37}\,{\rm erg}\left(\frac{l}{10^{4}\,\rm cm}\right)^{2}\left(\frac{\Delta R}{0.3\,\rm km}\right)\left(\frac{\epsilon}{10^{-3}}\right)^{2}.$
The strain energy is smaller than the magnetic energy released in the same area during the crack motion.
A global starquake may happen when the shear strain due to a spin-down-induced shape change exceeds a threshold. Note that a neutron star is thought to be a fluid star with a thin solid crust. The equilibrium configuration of a rotating neutron star leads to an ellipsoid rather than a perfect sphere, in which elastic energy is accumulated in the crust (Baym & Pines 1971). This global starquake is also associated with a glitch and consists of three steps: (1) the normal spin-down phase, which begins at the end of the last glitch and during which the elastic energy is accumulated; (2) the glitch epoch, during which the star loses its elastic energy; and (3) the glitch phase, during which the star changes its shape and sets up a new equilibrium at the end of this phase (Zhou et al. 2004; Peng & Xu 2008; Zhou et al. 2014). The total energy of the star is given by (Baym & Pines 1971)

### v2

@@ Abstract
With a growing sample of fast radio bursts (FRBs), we investigate the energy budget of different power sources within the framework of magnetar starquake triggering mechanism. During a starquake, the energy can be released in any form through strain, magnetic, rotational, and gravitational energies. The strain energy can be converted from other three kinds of energy during starquakes. The following findings are revealed: 1. The crust can store free magnetic energy of $\sim 10^{46}$ erg by existing toroidal fields, sustaining $10^{6}$ bursts with frequent starquakes occurring due to crustal instability. 2. The strain energy develops as a rigid object spins down, which can be released during a global starquake accompanied by a glitch. However, it takes a long time to accumulate enough strain energy via spindown. 3. The rotational energy of a magnetar with $P\lesssim 0.1\rm\,s$ can match the energy and luminosity budget of FRBs. 4. The budget of the total gravitational energy is high, but the mechanism and efficiency of converting this energy to radiation deserve further exploration.
@@ 1 Introduction
Fast radio bursts (FRBs) are millisecond-duration flashes, signifying extremely coherent radiation due to their extremely high bright temperatures. Although they are similar in some aspects to single pulses from radio pulsars, the energy budget of FRBs is much higher. As a result, the physical origin(s) of FRBs is still a mystery (Zhang 2023).
@@
A strong connection between magnetars and FRBs was established since the discovery of FRB 20200428D, an FRB-like burst from a Galactic magnetar, SGR J1935+2154 (Bochenek et al. 2020; CHIME/FRB Collaboration et al. 2020). Based on the distance from the emission site to the central engine, magnetar FRB models can be generally divided into two categories: emission within the magnetosphere (pulsar-like models, Katz 2014; Kumar et al. 2017; Yang & Zhang 2018; Wang et al. 2019; Lu et al. 2020; Cooper & Wijers 2021; Wang et al. 2022; Zhang 2022; Liu et al. 2023; Wang et al. 2024; Qu & Zhang 2024), and emission from a relativistic shock region far outside the magnetosphere (GRB-like models, Metzger et al. 2019; Beloborodov 2020; Margalit et al. 2020; Chen et al. 2022; Khangulyan et al. 2022). The observations of polarization measurements, including diverse polarization angle swings (Luo et al. 2020b) and high degrees of circular polarization (Jiang et al. 2022; Xu et al. 2022; Zhang et al. 2023b), favor the pulsar-like models (Qu & Zhang 2023), even though there are also some attempts to interpret these phenomena within the GRB-like models (Iwamoto et al. 2024).
@@
Starquakes from magnetars have been proposed as a promising triggering mechanism for FRBs (Wang et al. 2018; Suvorov & Kokkotas 2019; Wadiasingh & Timokhin 2019; Jiang et al. 2020; Li et al. 2022a). The FRB burst sequences share remarkably similar characteristics in energy distribution and temporal occurrence with Earthquakes (Du et al. 2023; Totani & Tsuzuki 2023; Wang et al. 2023; Tsuzuki et al. 2024). The quake model predicts that FRBs are associated with glitches and quasi-periodic oscillations (QPOs). Surprisingly, a giant glitch was measured days before FRB 20200428D (Mereghetti et al. 2020; Li et al. 2021a; Ridnaia et al. 2021), and later another glitch was found to be accompanied by three FRB-like radio bursts in subsequent days (Younes et al. 2023). A QPO of 40 Hz was reported from an X-ray burst during the active epoch of FRB 20200428D (Li et al. 2022b).
@@
Several FRB sources have been observed to repeat more than a thousand times (Li et al. 2021b; Xu et al. 2022; Zhang et al. 2023b), suggesting that an active central engine and a large energy reservoir are needed. The prospect of detecting more than 10 thousand bursts from an FRB source would pose significant challenges on the potential energy budget of a magnetar. In this paper, we investigate the energy budget and the mechanism that can be released, within the framework of the starquake model for a neutron star or a magnetar. We consider the energy budget of repeating FRBs in Section 2. We discuss the energy release during starquakes by invoking strain, magnetic, rotational, and gravitational energy in Section 3. The results are discussed in Section 4 and summarized in Section 5.
@@ 2 Energy budget of an FRB source
As a cosmological burst, an FRB can release a huge amount of energy. For a nominal Gpc distance $D$, the burst energy of an individual burst can be estimated as
$\delta E$ $\simeq 4\pi D^{2}f_{b}F_{\nu}\Delta\nu$ (1)
@@
where $f_{b}=\delta\Omega/(4\pi)$ is the beaming factor for an individual burst, $F_{\nu}$ is the received fluence and $\Delta\nu$ is the bandwidth of the telescope (for a narrow FRB spectrum). The emission solid angle $\delta\Omega$ could be as small as $1/(\gamma^{2})\sim(10^{-4}-10^{-6})$ due to relativistic beaming (Wang et al. 2022; Zhang 2023). Even though the energy of individual burst is smaller by a factor of $f_{b}$ than the isotropic value, there is a factor of $f_{b}^{-1}$ more FRBs that do not beam toward Earth. So the total energy budget of an FRB remains the same as estimated based on the observations (Li et al. 2021b; Xu et al. 2022; Zhou et al. 2022) because the two factors cancel out.
However, there might be a global emission beam factor. The bursts may not be emitted isotropically. The emission cone of individual bursts is confined to a global fan-beam which has a solid angle of $\Delta\Omega$ [^1] Conal regions from a magnetospheric rotator are considered as the global beaming scenario, but a more general geometry is possible., leading to a global beam factor of $F_{b}=\Delta\Omega/(4\pi)$, with $\Delta\Omega\gg\delta\Omega$. In this case, there is only a small fraction of bursts detected with altogether $\Delta\Omega/\delta\Omega$ bursts most of that are missed, so that the global energy budget can be smaller by a factor of $F_{b}$ from the isotropic value (Zhang 2023). Therefore, when considering the total burst energy/luminosity, the global beaming factor rather than the beaming factor of an individual burst should be adopted.
It is necessary to introduce the luminosity/energy function of FRBs when a large sample of FRBs is considered. Some statistical studies show that the luminosity/energy function of bursts from the same source can be characterized by a power law distribution, $N(E)\propto E^{-p}$ with indices ranging from $1-2$ (Li et al. 2021b; Xu et al. 2022; Zhou et al. 2022; Zhang et al. 2023b). For the FRB population involving different sources, the energy distribution can also be described by power laws with the similar index range (Cao et al. 2018; Luo et al. 2018; Lu & Piro 2019; Lu et al. 2020; Luo et al. 2020a). If the index is not strictly equal to 1 or 2, the average burst energy is calculated to
@@
Note that most bursts are missing due to the duty cycle. The total observed number and their energy budget of some actively repeating FRBs within duty cycle are summarized in Table 1. The physical duty cycle and burst rate evolution are complex. We assume that the repeater source has a lifetime $\tau_{\rm life}$ and consider the average bursting rate to be independent of energy. If the activity level remains unchanged during the lifetime, the total energy budget of the source over the lifetime can be estimated as (Zhang 2023)
@@
20121102A 2370 $4.1\times 10^{41}$ Spitler et al. (2016); Michilli et al. (2018); Zhang et al. (2018a);
@@ 3 Energy release during starquake
3.1 Strain energy from crustal cracking
Starquakes as a leading scenario to trigger FRBs can suddenly release a huge amount of energy. It has been discussed as the energy source for some high-energy phenomena since the elastic energy released in the crust excites oscillations and the induced electric field accelerates charges to stream outflow (Tsygan 1975; Fabian et al. 1976; Muslimov & Tsygan 1986; Epstein 1988; Blaes et al. 1989). The strain energy released in the crack region of $l^{2}$ is given by
$\delta E_{\rm str}$ $=2\pi l^{2}\Delta R\mu\epsilon^{2}$ (4)
$=4.7\times 10^{37}\,{\rm erg}\left(\frac{l}{10^{4}\,\rm cm}\right)^{2}\left(\frac{\Delta R}{0.3\,\rm km}\right)\left(\frac{\epsilon}{10^{-3}}\right)^{2},$
where $\epsilon$ is the strain of the crust and $\mu\approx 2.5\times 10^{30}\,\rm erg\,cm^{-3}$ is the shear modulus of the crust (Thompson & Duncan 1995). This energy release is consistent with the budget for an individual burst.
During the crustal motion, some strain energy is accumulated and it would be released when the stress exceeds a threshold. The starquakes may be induced via some magnetic activities or crust collapse. Also, starquakes may be associated with spindown glitches in which strain energy is accumulated via spindown-induced shape change. These processes are also accompanied by magnetic, rotational, and gravitational energy release, which are summarized in Figure 1. The possible processes involve a magnetar directly releasing these energies, partially converting these energies to strain energy. Crustal motion can generate seismic waves leading to oscillations of the star. The toroidal oscillation of the star enables the Goldreich-Julian density (Goldreich & Julian 1969) to deviate from the traditional value due to rotation, which provides additional voltage to accelerate charged particles to emit photons (Lin et al. 2015). The energy evolutionary paths are summarized in Figure 2. Some situations occur alternatively and some processes take place successively or simultaneously. We discuss the release mechanisms of these energies independently in the following.
Figure 1: A schematic cartoon for energy stored and released in a magnetar. Starquakes can cause the footpoints of magnetic field lines to move, leading to magnetic reconnection. Gravitational energy may be released if the bulk shrinks, and part of it may convert into strain energy in the crust, and some strain energy may be released directly. Rotational energy can be released via quakes and spindown, with the latter being the direct energy source for FRBs.
@@ 3.2 Magnetic energy released of starquake
Within the magnetar scenario, the magnetic energy can be released from the crust (and possibly even from the interior of the crust) into the magnetosphere, when the pressure induced by the internal magnetic field exceeds a threshold stress. The magnetic energy dominates the crustal deformation energy if the magnetic field strength $B>B_{\mu}=(4\pi\mu)^{1/2}=5.6\times 10^{15}\,\rm G$. Suppose the magnetic field is deformed away from magnetostatic equilibrium by an amount of $\delta B$. In that case, the elastic stress suddenly balances the Maxwell stress, i.e., $\mu\epsilon\simeq B\delta B/(4\pi)$. The largest yield strain of the crust is estimated as $\sim 10^{-4}-10^{-2}$. The effect of ohmic decay is neglected for young magnetars.
@@
where $B_{s}$ is the magnetic field strength on the stellar surface and $\Delta R$ is the crustal thickness. There is a height where charge density is insufficient to supply the current required by the wave and an electric field is formed to accelerate particle bunches, converting the magnetic energy into emissions.
The timescale of the energy release process in an area of $l^{2}$ is related to some instability growth. A displacement of the magnetic footpoints on the surface can be caused by the diffusion of the internal magnetic field, which is reminiscent of the turbulent convective motions in the Sun. The exchange in the positions of the footpoints is driven by the interchange instability (Thompson & Duncan 1993). We take the density of the neutron drip $\rho_{n}$ as the crustal density and then the Alfvén speed in the crust is written as $B_{s}/(4\pi\rho_{n})^{1/2}$. The growth time of the instability driven by the external magnetic field is
$t_{\rm ex}\sim l/v_{A}\lesssim 2.2\times 10^{-5}\,{\rm s}\left(\frac{l}{10^{4}\,\rm cm}\right)\left(\frac{B_{s}}{10^{15}\,\rm G}\right)^{-1}.$ (6)
@@
The solid crust of a magnetar plays a crucial role in various high-energy activities such as FRB. Elastic stress accumulates with the evolution of the internal $B$-field and stores a large amount of energy in the form of elastic stress. The total magnetic energy stored in the crust is
@@
where $R$ is the stellar radius. Starquakes can occur anywhere on the star’s surface in principle so that the global beam factor is close to an order of 1 even if it can be small. The burst number supported by the crustal magnetic energy is estimated by
@@
The magnetic energy stored in the magnetosphere is higher than that stored in the crust and can support $\sim 10^{5}$ bursts. However, it is difficult to release such magnetic energy in a dipolar configuration, because the dipole is the ground state of the magnetic field. The free energy that magnetic fields can release favors multipoles or twisted magnetic fields.
Some toroidal magnetic field components might anchor in a highly conducting crust of a neutron star (Thompson et al. 2002). The toroidal magnetic field can be transferred from the interior to the crust via crustal differential rotation along gravitational equipotential surfaces. A sudden crust quake can twist the magnetic field in the outer magnetosphere (Beloborodov 2009). The magnetic energy of the twisted field is then launched into the magnetosphere.
Considering that the toroidal magnetic field in the crust is $10^{16}$ G, the energy can support $\sim 10^{6}$ bursts based on Equation (9). The elastic strain can hardly balance the magnetic strain because of $B>B_{\mu}$ so that the crust is unstable and frequent starquakes may occur at the surface. In this case, the magnetic energy stored in the crust is at least $6.3\times 10^{46}$ erg by taking $B=B_{\mu}$ in Equation (5). A simulation has found that if the magnetic field has a strong toroidal component, the crust is most prone to rupture, and in this case, the epicenter of crustal earthquakes is near the equator (Lander et al. 2015). By assuming that the magnetic strain is comparable to gravity, there might be an upper limit of magnetic field of $\sim 10^{18}$ G, leading to $\sim 10^{10}$ bursts, which is a strict upper limit.
3.3 Rotational energy release via quake and spindown
We discuss the release of rotational energy via a global glitch and directly convert it into radiations.
3.3.1 Spindown-induced glitch with starquake
A global starquake may happen when the shear strain due to a spindown-induced shape change exceeds a threshold. Note that a neutron star is thought to be a fluid star with a thin solid crust. The equilibrium configuration of a rotating neutron star leads to an ellipsoid rather than a perfect sphere, in which elastic energy is accumulated in the crust (Baym & Pines 1971). This global starquake is also associated with a glitch and consists of three steps: (1) the normal spindown phase, which begins at the end of the last glitch and during which the elastic energy is accumulated; (2) the glitch epoch, during which the star loses its elastic energy; and (3) the glitch phase, during which the star changes its shape and sets up a new equilibrium at the end of this phase (Zhou et al. 2004; Peng & Xu 2008; Zhou et al. 2014). The total energy of the star is given by (Baym & Pines 1971)